[Brahms-l] Shift report 20050315 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue Mar 15 2005 - 08:00:53 EST
Shifter leader was SB
Shifters were 


1. Beam from previous shift continued until 4 am. Next store around 
6:20 am. Machine development to begin with current store around 9:30 am.

2.  Run No.  FS @6 deg-@B861   MRS @45 deg-@A1050
---------------------------------------------------
   13795      trig6(530k)        trig3(433k)
   13796      trig6(427k)        trig3(210k)
---------------------------------------------------
    Run No.  FS @8 deg-@B861  MRS @45 deg-@A1050
---------------------------------------------------
    13797      trig6(103k)       trig3(178k)

3. No detector problems observed

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Mar 15 08:01:21 2005

This archive was generated by hypermail 2.1.8 : Tue Mar 15 2005 - 08:01:30 EST